Solve using Cramers Rule xty - 2 =o 2 4x -y +22 =30 (2) 1x +-42 = -15 )
Answer #1

solution: Given system of equations x +y - 2 = 0 4X - Y +22=30 x + y - 42 = -15 Here the coefficient matrix is 1 1 1 -1 A = 1=> x = 0-30(-3) -15(1) 15 90 - 15 15 75 c 15 - 5 1 o 4 30 1 - 15 det 1 2 -4 A 1(-120+30)-4 (0-15)+1(8+30) 15 - 90 +60 +30 15Therefore the X= 5 , y = 0 solution is , Z = 5.